OpenStudy (anonymous):

Find the remainder when f(x) is divided by (x - k) f(x) = 4x^3 - 6x^2 + 3x + 1; k= -2 -3 -61 26 -49

OpenStudy (anonymous):

@Hero @Loser66 @robtobey

Nnesha (nnesha):

substitute k for -2 into the divisor (x-k) then solve for x

Nnesha (nnesha):

set it equal to zero \[ x-(\color{red}{-2})=0\] x= ?

Nnesha (nnesha):

and then substitute x for its value into f(x) function to find remainder

OpenStudy (anonymous):

\[\text{PolynomialQuotientRemainder}\left[4 x^3-6 x^2+3 x+1,x--2,x\right] \]{31 - 14 x + 4 x^2, -61} Quotient, Remainder

OpenStudy (anonymous):

Above from Mathematica v9

OpenStudy (anonymous):

x = -2?

Nnesha (nnesha):

yes right now substitute x for -2 \[f(x) = 4x^3 - 6x^2 + 3x + 1\] then solve right side

OpenStudy (anonymous):

x--2 = x + 2

OpenStudy (anonymous):

oh okay i see, it sums up to -61

Nnesha (nnesha):

yes right.

OpenStudy (anonymous):

okay thank you both

Nnesha (nnesha):

yw